OpenStudy (anonymous):

Solve the exponential equation. Express your solutions in exact form only. Please show all of your work. 2e^x=(√10) e^x +3 Solve the exponential equation. Express your solutions in exact form only. Please show all of your work. 2e^x=(√10) e^x +3 @Mathematics

OpenStudy (anonymous):

like solving \[2z=\sqrt{10}z+3\] then replace z by \[e^x\] and solve for x using logs

OpenStudy (anonymous):

2e^x=(√10) e^x +3 2e^x-(√10) e^x =3 e^x[2-(√10) ] =3 e^x=3/[2-(√10) ] taking ln on both sides x=ln{3/[2-(√10) ]}

OpenStudy (anonymous):

wait a second

OpenStudy (anonymous):

there is no solution to this because you have \[e^x=\frac{3}{2-\sqrt{10}}\] which is a negative number and \[e^x>0\] for all x